# ziji: a model to calculate X-ray spectra of black hole X-ray binaries

### v1

@@ 4. Discussion and conclusions
We note that the spectra presented in the previous section have been obtained assuming the disk electron density $n_{\rm e}=10^{22}$ cm-3, which is the maximum value allowed by the reflionx code. For lower values of the electron density, the relativistic reflection spectra do not present strong reflection features, as the ionization of the accretion disk is very high. This would not be consistent with observations, where we can see strong reflection features even in the X-ray spectra of Galactic black holes with Eddington-scaled luminosities of 0.3 and more. High density reflection models seem thus strictly necessary if we want to fit the spectra of black hole X-ray binaries (Liu et al., 2023).

### v2

@@ 4. Discussion and conclusions
We note that the spectra presented in the previous section have been obtained assuming the disk electron density $n_{\rm e}=10^{22}$ cm-3, which is the maximum value allowed by the reflionx code. For lower values of the electron density, the relativistic reflection spectra do not present strong reflection features, as the ionization of the accretion disk is very high. This would not be consistent with observations, where we can see strong reflection features even in the X-ray spectra of Galactic black holes with Eddington-scaled luminosities of 0.3 and more. As shown in Fig. 5, $\xi>10$ erg cm s-1 in the inner part of the accretion disk even for $\lambda=0.01$. If we employed the disk electron density $n_{\rm e}=10^{15}$ cm-3, which is still the most widely used value in current reflection studies, the disk would be fully ionized and the reflection spectrum would have no clear emission lines. This point is often overlooked in the literature because in current reflection models there is no link between the reflection spectrum and the Eddington luminosity of the source, but actually current reflection models would be suitable to fit only the data of black hole binaries with extremely low Eddington luminosity. High density reflection models seem thus strictly necessary if we want to fit the spectra of black hole X-ray binaries (Liu et al., 2023).
We also note that our model does not include the returning radiation in the calculation of the thermal spectrum of the disk. As shown in Li et al. (2005), the returning radiation has the effect to heat the accretion disk and the thermal spectrum calculated including the returning radiation is very similar to a thermal spectrum calculated without the returning radiation but with a higher value of the mass accretion rate. In the analysis of the spectra of a source we can measure an effective mass accretion rate, which is slightly higher than the actual mass accretion rate because the calculations do not include the effects of the returning radiation.
@@
If we compare the reflection spectra for $\rho=100$ and for $\rho=1$ in Fig. 4, we do not see clear differences. This means that the the contribution of the returning radiation of the thermal spectrum in the generation of the reflection spectrum is still subdominant when $\rho=1$. The reflection spectra with $\rho=0$ are instead clearly different from those with $\rho=1$. The role of the returning radiation of the thermal spectrum is certainly important when a source is in a soft state, but it can be presumably ignored during intermediate and hard states. A quantitative analysis on when the contribution of the returning radiation of the thermal spectrum is significant is beyond the scope of this work and does not depend only on the value of $\rho$. As shown in Fig. 6, even the coronal height $h$ can affect the strength of the Compton hump. The characteristics of the X-ray detectors and of the source are also quite important. The Compton hump region seems to be the most sensitive part of the spectrum and therefore an observation in which we can measure the Compton hump well is probably suitable to measure the parameter $\rho$ and distinguish a source in which the reflection spectrum is mainly produced by the illumination of the disk by a corona from a source in which the reflection spectrum is mainly generated by the returning radiation of the thermal component.